Consider a market with two firms in Cournot (quantity) competition. Market demand is given by q(p) = a − p. Each firm faces a constant marginal cost of c. a. (15 points) Suppose that the government imposes a unit tax of δ, so that if a firm sells q units of the good, that firm owes q · δ to the government. Find the equilibrium quantity, price paid by consumers, consumer surplus, and tax revenue. Your answers should be functions of a, τ , and c. Make sure you box your answers.
Answer #1

q (P) = a-P, Coumot dwopoly . Q = a-P, Q = 9+92 MC= 0, perunit tax=8, then New MC of firms = 8+c. then = (P-Mc)q, = (a-q4-92-Page No : 1P (at & 8 + 8c/ | Price paid SO Page No EqMO - (Q--C)/3 cs = $[99-9 [ a- (a +28+86)] CS = 1 a 8-9 ( 3a-a-28-20) =